OpenStudy (anonymous):

i need help graphing this o Produce at least 30 of product #1. x >= 30 shirt o Produce at least 20 of product #2. y >= 20 jean 3x + 5y <= 340 10x + 15y

OpenStudy (anonymous):

somebody.....anybody ?!

OpenStudy (anonymous):

im all aloneeeeee

jimthompson5910 (jim_thompson5910):

to graph x >= 30, you draw a vertical line at x = 30 and shade to the right

jimthompson5910 (jim_thompson5910):

to graph y >= 20, you draw a horizontal line at y = 20 and shade above the line

jimthompson5910 (jim_thompson5910):

to graph 3x + 5y <= 340, you graph the line 3x + 5y = 340 and then you shade below the line

jimthompson5910 (jim_thompson5910):

The last inequality is missing the inequality sign and the right sdie

jimthompson5910 (jim_thompson5910):

side*

OpenStudy (anonymous):

this all goes on 1 graph right?

jimthompson5910 (jim_thompson5910):

But anyways, the idea is to graph each inequality. Once you have all the inequalities graphed, you shade the region that's common to ALL inequalites

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

hold on let me draw it ! stay with me now ! dont leave me D:

jimthompson5910 (jim_thompson5910):

alright, go ahead

OpenStudy (anonymous):

okay i have the first 2 down ( i may sound retarded haha) but how do i graph 3x + 5y = 340?

jimthompson5910 (jim_thompson5910):

you're not retarded

jimthompson5910 (jim_thompson5910):

you just need more practice

jimthompson5910 (jim_thompson5910):

to graph any line, you need two points

jimthompson5910 (jim_thompson5910):

To get any point, you plug in any for x and solve for y. 3x + 5y = 340 3(0) + 5y = 340 .. plug in x = 0 0 + 5y = 340 5y = 340 y = 340/5 y = 68 So one point is (0,68)

jimthompson5910 (jim_thompson5910):

To find another point, we just use another x value, say x = 10 3x + 5y = 340 3(10) + 5y = 340 30 + 5y = 340 5y = 340 - 30 5y = 330 y = 330/5 y = 66

jimthompson5910 (jim_thompson5910):

So two points that are on 3x + 5y = 340 are (0,68) and (10,66)

jimthompson5910 (jim_thompson5910):

Plot those two points and draw a line through them to graph 3x + 5y = 340

OpenStudy (anonymous):

ohhhhh okay i get it now ! thank you sooooso much :)

jimthompson5910 (jim_thompson5910):

you're welcome